### v1

# Exact real time dynamics with free fermions in disguise

Abstract
We consider quantum spin chains with a hidden free fermionic structure, distinct from the Jordan-Wigner transformation and its generalizations. We express selected local operators with the hidden fermions. This way we can exactly solve the real time dynamics in various physical scenarios, including the computation of selected dynamical two point functions and Loschmidt amplitudes, in continuous or discrete time. In the latter case we build quantum circuits that can be implemented on a quantum computer. With this we extend the family of classically simulable quantum many body processes.
Introduction.— Free fermions play an important role in multiple areas of theoretical physics, due to their exact solvability and the simplicity of the computations with them. The two-dimensional classical Ising model is solvable by free fermions [1], just like many one dimensional quantum spin chains, such as the Ising chain or the XX model [2]. The free fermionic Kitaev chain [3] is an important candidate for fault tolerant quantum computing. In two dimensions the honeycomb lattice model is a free fermionic system that hosts anyons [4]. Free fermions are important also in quantum information theory: quantum circuits based on the so-called matchgates are free fermionic, and they enable classical simulability of quantum processes [5, 6, 7]. Free fermions also appear in the tensor network models of the holographic principle, where they lead to an efficient contraction of the tensor network [8].
Given their importance it is natural to ask: What is the widest class of physically meaningful models which can be solved by free fermions? How can we construct free fermionic operators in systems that are inherently bosonic? And what is the true computational advantage of the free fermions?
In quantum spin chains fermions can be constructed using the Jordan-Wigner (JW) transformation [9] (for generalizations see [10, 11, 12, 13, 14]). However, the JW transformation does not encompass all possibilities. In the last couple of years a number of models have been found which can be solved by hidden free fermionic structures [15, 16, 17, 18, 19, 20] (for earlier examples see [21, 22, 23]). These models typically involve 4-body or higher interactions when expressed using the JW fermions; but the Hamiltonian can be diagonalized by the hidden fermionic operators. These are highly non-local in the original spin operators, but also in the JW fermions. The solution of hidden-fermion models is also different from a generalized JW transformation in that it is not possible to express every term in the Hamiltonian as a bilinear in the hidden fermions [13, 14, 18]. Previous works focused on computing the spectrum and the ground state properties of such models [24, 25, 26, 27, 28] but the practical advantage of the free fermionic structures has not yet been demonstrated beyond the computation of the spectrum.
In this work we establish contact with local physics, focusing on the “free fermions in disguise” (FFD) model of Fendley [15]. We express selected local operators using the hidden fermions, thereby partially solving the “inverse problem”. Afterwards many dynamical scenarios become analytically tractable, demonstrating the usefulness of hidden free fermions.
Besides Hamiltonian dynamics we also consider discrete time evolution in quantum circuits with special geometries, compatible with the FFD model. These circuits can be realized in present day quantum computers. With this we demonstrate that free fermions in disguise lead to classically simulable quantum processes and this could be used for benchmarking quantum computers.
Model.— We consider one dimensional quantum spin chains and quantum circuits which are built on selected representations of an abstract algebra. We call it the FFD algebra. It is defined by the generators $h_{j}$, $j=1,2,\dots,M$ which satisfy $h_{j}^{2}=1$, $h_{j}=h_{j}^{\dagger}$ and the commutation relations
$\begin{split}\{h_{j},h_{j+1}\}=\{h_{j},h_{j+2}\}=0\\ [h_{j},h_{k}]=0,\quad|j-k|>2\,.\end{split}$ (1)
These relations are understood without periodicity in the indices, therefore the algebra always describes a system with open boundaries.
We consider a family of models defined by the Hamiltonians
$H=\sum_{j=1}^{M}b_{j}h_{j}$ (2)
where $b_{j}\in\mathbb{R}$ are arbitrary coupling constants. We specify a representation of the abstract algebra on a spin 1/2-chain of length $L=M$. The concrete operators are given by $h_{1}=X_{1}$, $h_{2}=Z_{1}X_{2}$, and
$h_{j}=Z_{j-2}Z_{j-1}X_{j},\quad j\geq 3\,.$ (3)
Here and below $X_{j}$, $Y_{j}$ and $Z_{j}$ stand for the Pauli matrices acting on site $j$ of the spin chain. Other representations were treated in [15, 29].
Introducing Jordan-Wigner fermions as
$\chi_{2k-1}=Z_{k}\prod_{j=1}^{k-1}X_{j},\qquad\chi_{2k}=Y_{k}\prod_{j=1}^{k-1}X_{j}\,,$ (4)
we see that the $h_{j}$ are quartic in the fermions for $j\geq 3$: $h_{j}=-\chi_{2j-4}\chi_{2j-3}\chi_{2j-1}\chi_{2j}$.
One might wonder whether the model is solvable by a different choice of Jordan-Wigner transformation. It was proven in [18] that this is not possible: there are no Majorana fermions $\tilde{\chi}_{a}$ such that every operator $h_{j}$ is expressed as a product of two Majorana fermions [13, 14]. However, the parameters of the model can be specified so that it becomes Jordan-Wigner solvable. For example, if we set $b_{3j}=0$ for $j=1,2,\dots[M/3]$, then we can introduce a relabeling for the remaining terms as $\tilde{h}_{2j-1}=h_{3j-2}$ and $\tilde{h}_{2j}=h_{3j-1}$ for $j=1,2,\dots[M/3]$ and we find that these operators satisfy
$\{\tilde{h}_{j},\tilde{h}_{j+1}\}=0,\quad[\tilde{h}_{j},\tilde{h}_{k}]=0\text{for }|j-k|>1\,.$ (5)
These relations describe the operators in the Hamiltonian in the quantum Ising chain with a transverse magnetic field and the model becomes Jordan-Wigner solvable [30, 10, 14, 13]. In this special case all our results below become identical to those obtained by the standard methods; detailed comparisons will be presented elsewhere. In the following we consider generic choices of the parameters $b_{j}$.
Solution.— The model possesses a family of extensive conserved charges. They are given by the logarithmic derivatives of a transfer matrix $T_{M}(u)$, which can be defined via the recursion relation
$T_{M}(u)=T_{M-1}(u)-ub_{M}h_{M}T_{M-3}(u)\,,$ (6)
with the initial conditions $T_{M}(u)=1$ for $M\leq 0$. The order of $T_{M}(u)$ in $u$ is $S=[M+2/3]$. The Hamiltonian is obtained as $H=-\partial_{u}T_{M}(u)|_{u=0}$. The transfer matrices satisfy $[T_{M}(u),T_{M}(v)]=0$, therefore they also commute with the Hamiltonian. Furthermore, they satisfy an inversion relation
$T_{M}(u)T_{M}(-u)=P_{M}(u^{2})\cdot\mathbb{1}\,.$ (7)
Here $P_{M}(x)$ is a polynomial of order $S$, which is given by a recursion relation analogous to (6)
$P_{M}(x)=P_{M-1}(x)-xb_{M}^{2}P_{M-3}(x)\,.$ (8)
These polynomials and their generalizations were studied also in the mathematical literature [31, 32]; the values $P_{M}(-1)$ for $b_{j}\equiv 1$ are known as Narayana’s cows sequence [33, 34].
It was proven in [15] that the model possesses a hidden free fermionic structure. There exist operators $\Psi_{k}$ and $\Psi_{-k}\equiv\Psi_{k}^{\dagger}$, $k=1,2,\dots S$, such that
$\{\Psi_{k},\Psi_{\ell}\}=\{\Psi_{-k},\Psi_{-\ell}\}=0,\quad\{\Psi_{k},\Psi_{-\ell}\}=\delta_{k,\ell}$ (9)
which diagonalize the Hamiltonian via
$H=\sum_{k=1}^{S}\varepsilon_{k}[\Psi_{k},\Psi_{-k}]\,.$ (10)
Here the $\varepsilon_{k}$ are interpreted as (half of the) single particle energies associated with the eigenmodes. They are given by $\varepsilon_{k}=(u_{k})^{-1}$, with $u_{k}$ being the roots of a polynomial $P_{M}(u^{2})$.
There are only $S=[(M+2)]/3$ fermionic eigenmodes for a size $M$ of the algebra. This implies that each level is degenerate with the same level of degeneracy which increases exponentially with $M$. The actual size of the degenerate sectors is $2^{L-S}$, which depends on the choice of representation, determining the volume $L$ for a given $M$. The smallest possible value for a faithful representation is $L=[(M+1)/2]$, which can be achieved for certain values of $M$ [29]. We find that even in these cases the degeneracy grows exponentially with $M$.
Fermions.— The construction of the fermionic operators is based on the so-called edge operator $\chi_{0}$. This is a Hermitian operator satisfying $\chi_{0}^{2}=1$, $\{\chi_{0},h_{1}\}=0$ and $[\chi_{0},h_{j}]=0$ for $j>1$. For certain values of $M$ we can choose $\chi_{0}$ to be a member of the FFD algebra [29], but it can also be an extra element. In the representation (3) we can conveniently choose $\chi_{0}=Z_{1}$.
The explicit formula for the fermionic operators is [15]
$\Psi_{\pm k}=\frac{1}{\mathcal{N}_{k}}T(\mp u_{k})\chi_{0}T(\pm u_{k}),$ (11)
where $\mathcal{N}_{k}$ is a known normalization factor, see  [^1] Supplemental Materials to “Exact real time dynamics with free fermions in disguise.”. The $\Psi_{k}$ depend on the choice of the auxiliary element $\chi_{0}$; this reflects a gauge freedom arising from the degeneracies [15, 29]. However, correlation functions of the elements of the FFD algebra are independent of this choice.
The inverse problem.— In order to compute correlation functions one needs to express local operators using the hidden fermions. In Jordan-Wigner solvable models with periodic boundary conditions the fermionic eigenmodes are Fourier transforms of local fermionic operators, and the inverse problem is solved by the inverse Fourier transform. However, the situation is more complicated in the FFD algebra. There are two main difficulties. First, the fermions do not have fixed operator numbers, when expressed using any set of local operators. Second, the FFD algebra has only $S$ fermionic eigenmodes for a given size $M$, therefore the inverse problem can not be fully solved, and we can only expect solutions for selected local operators.
Our strategy is the following. We express $\chi_{0}$ as a linear combination of the fermionic operators. We make the Ansatz
$\chi_{0}=\sum_{j=-S,j\neq 0}^{S}C_{j}{\Psi}_{j}$ (12)
and compute the $C_{j}$. Afterwards we construct a family of operators which are bilinears in the fermions.
Using the results of [15] we show in [35] that the expansion co-efficients are
$C_{j}=C_{-j}=\sqrt{\frac{P_{M\setminus 1}(u_{j}^{2})}{-u_{j}^{2}P_{M}^{\prime}(u_{j}^{2})}}\,,$ (13)
where $P_{M\setminus 1}(x)$ stands for a polynomial which is obtained analogous to (8) but with the substitution $b_{1}=0$ and the prime in $P_{M}^{\prime}$ denotes differentiation with respect to its argument.
It remains to be checked whether the expansion (12) is indeed complete. We performed numerical tests for small values of $M$ and discovered that for $M=3k$ and $M=3k+2$ the model possesses a Majorana zero mode as well. This is described by a Hermitian operator $\Psi_{0}$ with the properties $(\Psi_{0})^{2}=\mathbb{1}$, $[H,\Psi_{0}]=0$ and $\{\Psi_{0},\Psi_{k}\}=\{\Psi_{0},\Psi_{-k}\}=0$ for $k=1,2,\dots S$.
If we add the zero mode as
$\chi_{0}=\sum_{j=-S}^{S}C_{j}{\Psi}_{j},\quad C_{0}=\begin{cases}0&M\in 3\mathbb{N}+1\,,\\ \prod_{k=1}^{S}\frac{u_{k}}{\hat{u}_{k}}&\text{else},\end{cases}$ (14)
where $\hat{u}_{k}>0$ are the roots $P_{M\setminus 1}(\hat{u}_{k}^{2})=0$, then numerical tests and an analytic proof presented in [35] show that the expansion is indeed complete, and we obtained the solution of the inverse problem for $\chi_{0}$.
We can now express other operators using the fermions, by employing the FFD algebra together with (10). A simple way is to construct a series of operators $o_{j}$ via the recursion $o_{j}=[H,o_{j-1}]/2$, together with the initial condition $o_{0}=\chi_{0}$. This can be seen as a Krylov basis in operator space. $H$ is bilinear in the fermions, while $o_{0}=\chi_{0}$ is linear, therefore every $o_{j}$ is also linear in the fermions. In the FFD algebra representation every $o_{j}$ includes a factor of $\chi_{0}$. Having found the $o_{j}$, we construct various products $o_{j}o_{k}$ and take their linear combinations, to obtain a family of local operators that are bi-linear in the fermions. It is crucial that $\chi_{0}$ drops out from every such product.
For the first element we find $o_{1}=b_{1}h_{1}\chi_{0}$. Taking the product $o_{1}\chi_{0}=b_{1}h_{1}$ we compute $h_{1}$ as a bi-linear expression in the fermions. It follows from (10), (12), and the fermionic algebra that
$h_{1}=b_{1}^{-1}\sum_{j,k=-S}^{S}\varepsilon_{j}C_{j}C_{k}\Psi_{j}\Psi_{k}\,.$ (15)
For the next element in the Krylov basis we find $o_{2}=((b_{2}h_{2}+b_{3}h_{3})b_{1}h_{1}+b_{1}^{2})\chi_{0}$, and using this result we can eventually express the combination $b_{2}h_{2}+b_{3}h_{3}$ as a bi-linear in the fermions. However, we did not find a way to express $h_{2}$ and/or $h_{3}$ individually using the fermions. Afterwards the next simplest operator for which a bi-linear expression is found is a linear combination of $h_{4}$, $h_{5}$ and $h_{2}h_{3}h_{5}$ [35]. Our method expresses more and more complicated local operators using the fermions. At present we do not have a full classification as to which local operator is bi-linear, or perhaps of higher order in the hidden fermions.
Dynamics.— We compute infinite temperature two-point functions of operators localized around the boundary. Our main example is
$D(t)=\left\langle h_{1}(t)h_{1}(0)\right\rangle\equiv\text{Tr}(h_{1}e^{-iHt}h_{1}e^{iHt})/\text{Tr}(\mathbb{1})\,.$ (16)
This correlation function can be expressed as
$D(t)=\frac{1}{4b_{1}^{2}}\left((\dot{B}(t))^{2}-B(t)\ddot{B}(t)\right),\quad B(t)=\left\langle\chi_{0}(t)\chi_{0}(0)\right\rangle\,.$ (17)
The time evolution of the fermionic modes is given by $\Psi_{\pm k}(t)=e^{\pm 2i\varepsilon_{k}t}\Psi_{\pm k}$, and a standard computation yields
$\begin{split}B(t)&=\sum_{k=0}^{S}C_{k}^{2}\cos(\theta_{k})\,,\\ D(t)&=\frac{1}{4b_{1}^{2}}\sum_{k,\ell=0}^{S}C_{k}^{2}C_{\ell}^{2}\sum_{\sigma=\pm}(\epsilon_{k}-\sigma\epsilon_{\ell})^{2}\cos(\theta_{k}+\sigma\theta_{\ell}),\end{split}$ (18)
where $\theta_{k}=2\varepsilon_{k}t$.
Results.— The above formula can be evaluated in polynomial time for every $M$ and arbitrary set of coupling constants. It is natural to consider parameters with a period 3 staggering, in which case the model displays a rich phase diagram [15]. We numerically evaluated the correlation function $D(t)$ for different values of $M$ and various choices of the staggered parameters. We found that the correlation function can show various types of behaviour: it can decay to zero with some power law, it can converge to a non-zero value, and it can also show persistent oscillations around a non-zero value. Similar variation of dynamical boundary correlations was observed earlier in standard free fermionic systems [36, 37, 38]; a full analysis of the behaviour of $D(t)$ will be presented elsewhere.
Here we focus on the completely homogeneous case with $b_{k}=1$ for $k=1,\dots,M$. This is a multi-critical point in the phase diagram, where the energy gap (determined by the smallest $\varepsilon_{k}$) scales as $M^{-z}$ with the unusual exponent $z=3/2$. We computed a closed form result for $B(t)$ in the $M\to\infty$ limit, by transforming the finite sum into a contour integral. Our final result reads
$B(t)=\int_{0}^{\pi}\mathrm{d}p\ C^{2}(p)\cos(2t\epsilon(p)),$ (19)
where $p$ is a momentum-like variable,
$\epsilon^{2}(p)=\frac{\sin^{3}p}{\sin\left(p/3\right)\sin^{2}\left(2p/3\right)},\quad p\in[0,\pi]$ (20)
and $C^{2}(p)$ is a known function [35]. We also showed that $B(t)$ can be expressed alternatively using the generalized hypergeometric function, see [35]. The correlation function $D(t)$ is computed afterwards via (17).
We computed the large time asymptotics of $D(t)$, and found
$D(t)\approx\alpha\frac{\sin\left(3\sqrt{3}t+3\pi/4\right)}{t^{13/6}}\,,$ (21)
where $\alpha$ is a known constant [35]. The unusual scaling exponent $13/6$ arises from the combination of two asymptotic contributions to $B(t)$, which belong to universality classes with dynamical exponents $z=1$ and $z=3/2$ [35].
In Figure 1 we plotted the function $D(t)$ for finite $M$ and also in the $M\to\infty$ limit, together with the asymptotic form given above.
Figure 1: The boundary correlation function $D(t)$ in the thermodynamic limit $M\to\infty$, plotted against the asymptotic formula (21). The inset also shows a comparison with data from finite $M$.
Quantum circuits.— Let us now return to finite $M$ and consider discrete time evolution in quantum circuits. Our aim is to derive circuits that can be realized in present day quantum computers, and which are compatible with the hidden free fermions [29].
The transfer matrix $T_{M}(u)$ admits a factorization, which yields a unitary quantum circuit if $u$ is purely imaginary [15, 29]. This way we obtain a unitary matrix $\mathcal{V}(\delta)$ parametrized by $\delta\in\mathbb{R}$, given by
$\mathcal{V}(\delta)=G\cdot G^{T},\quad G=u_{1}u_{2}\cdots u_{M-1}u_{M},$ (22)
where $u_{j}=e^{i\varphi_{j}h_{j}}$ are local unitary gates with the angles $\varphi_{j}$ given by the recursion
$\tan(2\varphi_{j})=-\delta b_{j}\cos(2\varphi_{j-1})\cos(2\varphi_{j-2})$ (23)
together with $\varphi_{0}=\varphi_{-1}=0$. For small values of $\delta$ we get $\mathcal{V}(\delta)\approx 1-i\delta H$, therefore $\mathcal{V}(\delta)$ can be seen as a “free fermionic Trotterization” of the FFD Hamiltonian, with $\delta$ playing the role of the discretization time step.
We choose the circuit $\mathcal{V}(\delta)$ to generate discrete time evolution. We define dynamical two-point functions as
$\mathcal{D}(\delta,N)=\text{Tr}(h_{1}\mathcal{V}(\delta)^{N}h_{1}\mathcal{V}(-\delta)^{N})/\text{Tr}(\mathbb{1})\,.$ (24)
[…]
In the representation (3) the local gates $u_{j}$ are controlled unitaries in the $Z$-basis:
$u_{j}=\frac{1+Z_{j-2}Z_{j-1}}{2}e^{i\varphi_{j}X_{j}}+\frac{1-Z_{j-2}Z_{j-1}}{2}e^{-i\varphi_{j}X_{j}}\,.$ (25)
These gates can be decomposed into 4 CNOT gates and 1 single qubit rotation [39, 35]. Therefore the circuit $\mathcal{V}(t)$ can be implemented on a quantum computer.
Quantum computation.— Our results for $\mathcal{D}(\delta,N)$ above can be seen as classical simulability of the Heisenberg time evolution of a selected operator. Now we derive a new quantum circuit, where the same quantity $\mathcal{D}(\delta,N)$ describes the fidelity in a certain quantum computation protocol. Therefore, we find classical simulability of the evolution of quantum states, in a special quantum circuit.
The idea is to transform the Heisenberg operator evolution into dynamics of quantum states. This is achieved by an operator-state correspondence in the FFD algebra. The elements of the FFD algebra can always be normal ordered; we represent any operator product as
$h_{a_{1}}h_{a_{2}}\dots h_{a_{n}},\qquad a_{1}<a_{2}<\dots<a_{n}\,.$ (26)
Such an operator can be represented by a bit string of length $M$, by writing down zeroes or ones depending on whether a specific $h_{j}$ is included in the product or not. If the zeroes and ones are interpreted as up and down spins, then we map the elements of the FFD algebra to the computational basis of an auxiliary spin-1/2 chain with length $M$.
Now we represent the Heisenberg time evolution of operators under the circuit $\mathcal{V}(\delta)$ as a new quantum circuit on the auxiliary chain. We use the factorization of $\mathcal{V}(\delta)$ into the local gates, and compute the conjugations
$e^{-i\varphi_{j}h_{j}}[\dots]e^{i\varphi_{j}h_{j}}\,,$ (27)
where []delimited-[][$\dots$][ … ] stands for any operator product of the form (26). The operator $h_{j}$ commutes or anticommutes with the operator product if $Z_{j-2}Z_{j-1}Z_{j+1}Z_{j+2}=\pm 1$ on the auxiliary chain, respectively. Only the second case yields a non-trivial action. A careful computation shows that such a conjugation step is represented by a controlled unitary acting on the auxiliary chain:
$U_{j}=\Pi^{(a)}+\Pi^{(b)}e^{2i\varphi_{j}X_{j}}+\Pi^{(c)}e^{-2i\varphi_{j}X_{j}}$ (28)
Here $\Pi^{(a,b,c)}$ are mutually orthogonal and complete projectors acting on the four sites $j-2,j-1,j+1$ and $j+2$. Their explicit form is given in [35], together with their factorization into 8 CNOT gates and 2 single qubit rotations.
Finally, the full circuit is given by
$V=G\cdot G^{T},\text{ where now }G=U_{1}U_{2}\cdots U_{M-1}U_{M}\,.$ (29)
We need to transform two-point functions like (24) to a quantity on the auxiliary chain. In the FFD algebra every non-identical operator product is traceless [29], so the trace selects the identity component. Every initial operator is represented by a computational basis state, the Heisenberg time evolution is represented by the circuit $V$, and taking the product with an initial operator (together with taking the trace) corresponds to taking a scalar product with a “final” computational basis state. For example, the initial operator $h_{1}$ corresponds to the initial state $|\ket{\psi_{0}}=\ket{\downarrow\uparrow\uparrow\uparrow\dots}$, while other operators bi-linear in the fermions correspond to linear combinations of basis states [35]. Finally, the two-point function (24) corresponds to
$\mathcal{D}(\delta,N)=\bra{\psi_{0}}V^{N}\ket{\psi_{0}}\,.$ (30)
This can be interpreted as the fidelity, or alternatively, as a Loschmidt amplitude in a “local quench”.
The key advantage of this formulation is that now $\mathcal{D}(\delta,N)$ can be measured in a quantum computer. The initial states (including those that are linear combinations of basis states [35]) can be prepared by shallow circuits, and the fidelity can be obtained for example by the SWAP-test [40, 41].
Discussion.— We used the hidden free fermionic structure to compute the time evolution of selected physical observables. All the problems considered here can be solved in polynomial time in $M$, including finding the roots $u_{k}$ and evaluating the summations for the obersvables. This means that we can classically simulate the corresponding quatum many body processes.
Our framework should be compared to circuits made up of matchgates [5, 6, 7], which are two-site gates whose logarithm is bi-linear in the Jordan-Wigner fermions. It follows that any product of the matchgates is still free fermionic; the only requirement is that each gate should act on two neighbouring sites.
By contrast, in our framework the gates in the circuit need to be arranged in a special way to preserve the hidden free fermionic property. In this work we treated only one arrangement, given by (29). Other arrangements were explored in [29]. Currently it is not known what is the most general class of circuits compatible with the free fermions, and this is a direction for future research.
There are differences also regarding the initial states. In the case of the matchgates the analytic computation is feasible for any basis state as initial state [5, 6, 7]. In our case the allowed states originate from those operators in the FFD algebra that can be represented by a combination of the fermionic operators. We presented examples of such initial states, but currently we don’t have a full characterization of the allowed states. We hope to return to this question in future work.
Acknowledgements.
We are thankful to János Asbóth, Paul Fendley, Esperanza Lopez, Lorenzo Piroli, Roberto Ruiz, German Sierra, Alejandro Sopena, Eric Vernier for stimulating discussions. We also benefitted from the workshop “Exactly Solved Models and Quantum Computing” organized at the Lorentz Center, Leiden, The Netherlands, in March 2024. The authors were supported by the Hungarian National Research, Development and Innovation Office, NKFIH Grant No. K-145904 and B.P. was supported by the NKFIH excellence grant TKP2021_NKTA_64.